Suppose the following demand and supply function:

Qd = 750 – 25P

Qs = -300 + 20 P


       i.           Find equilibrium price and quantity

     ii.           Find consumer and producer surplus  



1
Expert's answer
2021-06-02T09:38:21-0400

i. In equilibrium Qd = Qs, so:

750 – 25P = -300 + 20P,

45P = 1,050,

P = 23.33,

Q = -300 + 20*23.33 = 166.67 units.

ii. Consumer surplus is:

"CS = 0.5*166.67*(30 - 23.33) = 555.85."

Producer surplus is:  

"PS = 0.5*166.67*(23.33 - 15) = 694.18."